# Quasiparticle band picture bridging topology and strong correlations across energy scales

### v1

@@ I Introduction
Topological phases in solid-state systems have been widely explored for periodic models of noninteracting electrons, where physical observables are related to topological invariants that can be directly computed from the eigenstates of single-particle Hamiltonians [22, 23, 24], which are easy to diagonalize. However, this framework breaks down in the presence of electronic correlations, when the size of the Hamiltonian matrix scales with the Hilbert space rather than the system volume. In weakly correlated systems, effective single-particle descriptions, such as Density Functional Theory (DFT) [25, 26, 27], can offer invaluable insights for guiding material design [28, 29, 30], but they can fall short when interaction effects dominate over the band energy gain. Conversely, state-of-the-art methods for strongly correlated systems, such as Dynamical Mean-Field Theory (DMFT) [4, 31, 32], excel at describing local correlation effects, but lack a natural and directly accessible band-structure picture applicable at both low and high energy scales, which is important to describe the inter-band coherence effects relevant to topology. Moreover, DMFT and similar many-body methods can be computationally demanding across wide parameter regimes, critically hindering the extensive exploration of potential material candidates. This limitation is particularly relevant in the context of twisted material platforms, where the high degree of tunability underscores the richness of the phase diagram [33, 34, 35].
@@ II.1 Interacting Bernevig-Hughes-Zhang model
where $\Psi^{\dagger}_{\mathbf{k}}=\big{(}s^{\dagger}_{\mathbf{k}\uparrow},s^{\dagger}_{\mathbf{k}\downarrow},p^{\dagger}_{\mathbf{k}\uparrow},p^{\dagger}_{\mathbf{k}\downarrow}\big{)}$ are four component spinors and $H_{\mathrm{BHZ}}(\mathbf{k})$ is the $4\times 4$ matrix
@@
The phase diagram of the model in presence of a Coulomb repulsion between the electrons has attracted a lot of interest in the field [47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57]. The onsite Coulomb interaction projected onto the Wannier orbital basis can be written as
@@ II.3 Spectral representations of correlated electrons
is therefore also sometimes referred to as quasiparticle Hamiltonian within the scope of Fermi-liquid theory. As we will discuss in the following section, $H_{*}(\mathbf{k})=H_{*}(\epsilon\to 0,\mathbf{k})$ is commonly employed in the study of strongly correlated systems as a renormalized one-body effective Hamiltonian to characterize the integrated topology encoded in the Green’s function winding number [78, 79, 60]. To avoid confusion with $H_{\mathrm{qp}}$ in the gGut Ansatz, here we will use the term Fermi-liquid quasiparticle Hamiltonian for $H_{*}$.
@@ II.4 Topological invariants within ghost Gutzwiller
We are finally ready to examine how the band-structure emerging from $H_{\mathrm{qp}}(\mathbf{k})$ can be used to extract the topological invariants of the system. Specifically, following the previous discussion on spectral representations of correlated electrons, we demonstrate the physical relevance of computing Chern numbers for these quasiparticle bands, and their relation to the Chern number of the fully interacting system. We begin by revisiting a foundational tool in the study of correlated topological insulators (TIs): the single-particle Green’s function winding number [78, 79, 60]
@@
where $G(i\epsilon,\mathbf{k})$ is the interacting Green’s function in the Matsubara frequency domain. When perturbation theory holds, $W[G]$ coincides with the zero-temperature Hall conductance in units of $e^{2}/2\pi\hbar$ [78, 79, 60]. This makes the winding number a particularly powerful tool for methods that grant access to the full single-particle Green’s function, such as DMFT and gGut. However, the direct evaluation of (20) is often computationally expensive. A common alternative is to exploit the fact that this winding number is equivalent to the Chern number of the Hermitian matrix $-G(0,\mathbf{k})^{-1}$ [80], or equivalently $-G_{*}(0,\mathbf{k})^{-1}$ [60], which is related to the Fermi-liquid quasiparticle Hamiltonian in Eq. (11).
@@
especially if the edge modes of $H_{\text{qp}}(\mathbf{k})$ appear also in the spectrum of $G^{\text{gGut}}(i\epsilon,\mathbf{k})$. If the conjecture (25) is true, as we assume herein, then calculating $W\left[G^{\text{gGut}}\right]$ becomes considerably simplified: one merely requires to evaluate the Chern number of the quasiparticle bands.
We conclude by noticing that $H_{\mathrm{qp}}(\mathbf{k})$ may provide even more topological information than merely the winding number (20). The latter is, in fact, an integrated quantity that does not convey information about the energy and momentum resolved topological features useful for interpreting experimental spectra. Conversely, the quasiparticle Hamiltonian $H_{\mathrm{qp}}(\mathbf{k})$ in gGut offers this additional information without any additional computational cost. This feature grants access to a band structure that, unlike DFT, also encompasses the Hubbard bands, and thus enables a direct comparison with experimental data under the assumption that the quasiparticle bands accurately represent the genuine single-particle spectrum.
@@ V Topology of the Hubbard bands from topology of the ghost states
As highlighted in the approximate analytic argument, this shift is a genuinely nonperturbative feature of the high-energy spectrum in the Mott insulating regime, and is fully captured within the gGut framework. It is worth noting that a topological classification of MIs has also been proposed based on the structure of the zeros of the Green’s function: when nonlocal dynamical correlations are included—beyond what is accessible in single-site DMFT or gGut—dispersive zeros emerge within the Mott gap. These zeros can acquire a topologically nontrivial structure, contributing to a finite winding number of the Green’s function itself [54, 60]. Importantly, the existence of zeros within the Mott gap reflects the underlying structure of the Green’s function as a sum of contributions from the LHB and UHB poles; the presence of a pair of poles generically implies a zero in between. This observation suggests that, at least at leading order, the topology of the Hubbard bands and the topology of Green’s function zeros may originate from a common nonperturbative mechanism. Further exploration of this potential link could shed light on how local and nonlocal correlations cooperate in defining the topological structure of interacting systems. A detailed analysis will require cluster extensions of gGut that incorporate nonlocal correlation effects, which we leave for future studies.
@@ VII Conclusion
Overall, our study paints a convincing portrait of the gGut framework as a reliable, efficient and, most importantly, interpretable tool to study the interplay between correlation and topology. By restoring a quasiparticle perspective within a correlated setting, it brings the description of many-body phenomena closer to the language of band structure theory, unlocking the potential for predictive modeling in correlated materials. This capability paves the way to extend powerful paradigms, such as topological quantum chemistry [28, 29, 30], to strongly interacting regimes, beyond the traditional domain of weakly correlated systems. Moreover, it opens the door to high-throughput computational searches of correlated topological materials, accelerating the discovery of novel platforms for low-power electronics, spintronics, and quantum memory devices.
@@ Acknowledgments
We acknowledge insightful discussions with A. Amaricci, M. Capone, A. Marrazzo, R. Resta and J. Skolimowski. I.P. also gratefully acknowledges S. Giuli for valuable assistance with the numerical implementation during the initial stages of the project. G.B. is financially supported by the National Recovery and Resilience Plan PNRR MUR Project No. CN00000013-ICSC and by MUR via the PRIN 2020 (Prot. 2020JLZ52N-002) and PRIN 2022 (Prot. 20228YCYY7) programs. B.M. acknowledges support from a UKRI Future Leaders Fellowship [MR/V023926/1] and from the Gianna Angelopoulos Programme for Science, Technology, and Innovation.

### v2

@@ I Introduction
Topological phases in solid-state systems have been widely explored for periodic models of noninteracting electrons, where physical observables are related to topological invariants that can be directly computed from the eigenstates of single-particle Hamiltonians [22, 23, 24], which are easy to diagonalize. However, this framework breaks down in the presence of electronic correlations, when the size of the Hamiltonian matrix scales with the Hilbert space rather than the system volume. In weakly correlated systems, effective single-particle descriptions, such as Density Functional Theory (DFT) [25, 26, 27], can offer invaluable insights for guiding material design [28, 29, 30, 31, 32], but they can fall short when interaction effects dominate over the band energy gain. Conversely, state-of-the-art methods for strongly correlated systems, such as Dynamical Mean-Field Theory (DMFT) [4, 33, 34], excel at describing local correlation effects, but lack a natural and directly accessible band-structure picture applicable at both low and high energy scales, which is important to describe the inter-band coherence effects relevant to topology. Moreover, DMFT and similar many-body methods can be computationally demanding across wide parameter regimes, critically hindering the extensive exploration of potential material candidates. This limitation is particularly relevant in the context of twisted material platforms, where the high degree of tunability underscores the richness of the phase diagram [35, 36, 37].
@@ II.1 Interacting Bernevig-Hughes-Zhang model
where $\Psi^{\dagger}_{\mathbf{k}}=\big{(}s^{\dagger}_{\mathbf{k}\uparrow},s^{\dagger}_{\mathbf{k}\downarrow},p^{\dagger}_{\mathbf{k}\uparrow},p^{\dagger}_{\mathbf{k}\downarrow}\big{)}$ are four component spinors and $H_{\mathrm{BHZ}}(\mathbf{k})$ is the $\times 4$ 4 matrix
@@
The phase diagram of the model in presence of a Coulomb repulsion between the electrons has attracted a lot of interest in the field [49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59].
The onsite Coulomb interaction projected onto the Wannier orbital basis can be written as
@@ II.3 Spectral representations of correlated electrons
is therefore also sometimes referred to as quasiparticle Hamiltonian within the scope of Fermi-liquid theory. As we will discuss in the following section, $H_{*}(\mathbf{k})=H_{*}(\epsilon\to 0,\mathbf{k})$ is commonly employed in the study of strongly correlated systems as a renormalized one-body effective Hamiltonian to characterize the integrated topology encoded in the Green’s function winding number [80, 81, 62, 82]. To avoid confusion with $H_{\mathrm{qp}}$ in the gGut Ansatz, here we will use the term Fermi-liquid quasiparticle Hamiltonian for $H_{*}$.
@@
We note that pole-expansion based expressions have been explored in the context of correlated topology [83, 84, 85, 86]. These typically are constructed as few-pole approximations to an otherwise computed self-energy, e.g. through DMFT. In contrast, we stress that the pole-expansion of the Green’s function in gGut, and thus the quasiparticle band description, is naturally built into the Ansatz: the position and bandwidths of the ghost bands are fixed through the embedding approximation to the variational principle.
@@ II.4 Topological invariants within ghost Gutzwiller
We are finally ready to examine how the band-structure emerging from $H_{\mathrm{qp}}(\mathbf{k})$ can be used to extract the topological invariants of the system. Specifically, following the previous discussion on spectral representations of correlated electrons, we demonstrate the physical relevance of computing Chern numbers for these quasiparticle bands, and their relation to the Chern number of the fully interacting system.
To this end, we note that the gGut Green’s function introduced in Eq. (12) embodies the notion that the quasiparticle Hamiltonian, although optimized to describe ground state properties, also provides a meaningful approximation to single-particle excitations. This perspective becomes even more compelling when considering the robust quantization of topological invariants of the ground state. The energy- and momentum-resolved features of the approximate single-particle excitations in $H_{\text{qp}}(\mathbf{k})$ are expected to be protected by the intrinsic robustness of the topological invariants they support. In this sense, it is reasonable to expect that $H_{\text{qp}}(\mathbf{k})$ in Eq. (6) should properly describe the topological features of the original, interacting system. It is instructive to work out this description in a particular example, and thus the rest of this subsection is devoted to explicitly establishing the connection between the gGut quasiparticle Hamiltonian and the winding number of the physical Green’s function, a central quantity in the description of ground state topology in self-energy based methods like DMFT.
The single-particle Green’s function winding number can be written as [80, 81, 62]
@@
where $G(i\epsilon,\mathbf{k})$ is the interacting Green’s function in the Matsubara frequency domain. When perturbation theory holds, $W[G]$ coincides with the zero-temperature Hall conductance in units of $e^{2}/2\pi\hbar$ [80, 81, 62, 82]. This makes the winding number a particularly powerful tool for methods that grant access to the full single-particle Green’s function, such as DMFT and gGut. However, the direct evaluation of (20) is often computationally expensive. A common alternative is to exploit the fact that this winding number is equivalent to the Chern number of the Hermitian matrix $-G(0,\mathbf{k})^{-1}$ [87], or equivalently $-G_{*}(0,\mathbf{k})^{-1}$ [62], which is related to the Fermi-liquid quasiparticle Hamiltonian in Eq. (11).
@@
especially if the edge modes of $H_{\text{qp}}(\mathbf{k})$ appear also in the spectrum of $G^{\text{gGut}}(i\epsilon,\mathbf{k})$. If the conjecture (25) is true, as we assume herein, then calculating $W\left[G^{\text{gGut}}\right]$ does not require the explicit computation of the Green’s function: one merely requires to evaluate the Chern number of the quasiparticle bands.
We conclude by stressing that $H_{\mathrm{qp}}(\mathbf{k})$ can provide even more topological information than merely the winding number (20). The latter is, in fact, an integrated quantity that does not convey information about the energy and momentum resolved topological features useful for interpreting experimental spectra. Conversely, the quasiparticle Hamiltonian $H_{\mathrm{qp}}(\mathbf{k})$ in gGut offers this information as well without any additional computational cost. This feature grants access to a band structure that, unlike DFT, also encompasses the Hubbard bands, and thus enables a direct comparison with experimental data under the assumption that the quasiparticle bands accurately represent the genuine single-particle spectrum.
@@ V Topology of the Hubbard bands from topology of the ghost states
As highlighted in the approximate analytic argument, this shift is a genuinely nonperturbative feature of the high-energy spectrum in the Mott insulating regime, and is fully captured within the gGut framework. It is worth noting that a topological classification of MIs has also been proposed based on the structure of the zeros of the Green’s function: when nonlocal dynamical correlations are included—beyond what is accessible in single-site DMFT or gGut—dispersive zeros emerge within the Mott gap. These zeros can acquire a topologically nontrivial structure, contributing to a finite winding number of the Green’s function itself [56, 62, 62, 95, 82, 96]. Importantly, the existence of zeros within the Mott gap reflects the underlying structure of the Green’s function as a sum of contributions from the LHB and UHB poles; the presence of a pair of poles generically implies a zero in between. This observation suggests that, at least at leading order, the topology of the Hubbard bands and the topology of Green’s function zeros may originate from a common nonperturbative mechanism. Further exploration of this potential link could shed light on how local and nonlocal correlations cooperate in defining the topological structure of interacting systems. A detailed analysis will require cluster extensions of gGut that incorporate nonlocal correlation effects, which we leave for future studies.
@@ VII Conclusion
Overall, our study paints a convincing portrait of the gGut framework as a reliable, efficient and, most importantly, interpretable tool to study the interplay between correlation and topology. By restoring a quasiparticle perspective within a correlated setting, it brings the description of many-body phenomena closer to the language of band structure theory, unlocking the potential for predictive modeling in correlated materials. This capability paves the way to extend powerful paradigms, such as topological quantum chemistry [28, 29, 30, 31, 32], to strongly interacting regimes, beyond the traditional domain of weakly correlated systems. Moreover, it opens the door to high-throughput computational searches of correlated topological materials, accelerating the discovery of novel platforms for low-power electronics, spintronics, and quantum memory devices.
@@ Acknowledgments
We acknowledge insightful discussions with A. Amaricci, M. Capone, N. Lanatà, A. Marrazzo, R. Resta and J. Skolimowski. I.P. also gratefully acknowledges S. Giuli for valuable assistance with the numerical implementation during the initial stages of the project. G.B. is financially supported by the National Recovery and Resilience Plan PNRR MUR Project No. CN00000013-ICSC and by MUR via the PRIN 2020 (Prot. 2020JLZ52N-002) and PRIN 2022 (Prot. 20228YCYY7) programs. B.M. acknowledges support from a UKRI Future Leaders Fellowship [MR/V023926/1] and from the Gianna Angelopoulos Programme for Science, Technology, and Innovation.